OpenStudy (anonymous):

Find the solution to the system of equations using either substitution or the linear combination method. Show all your work. I will post the problem right after this is posted.

OpenStudy (anonymous):



OpenStudy (anonymous):

Ok so substitution is done by isolating 1 variable in 1 equation, so pick 1 equation and 1 variable then solving for that variable

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

i dont think thats right

OpenStudy (trojanpoem):

y = -34 - 2x 3x + 5y = -9 3x - 170 - 10x = - 9 -7x = 161 x = 161/-7 = -23 y = -34 + 46 = 12